OpenStudy (anonymous):

Trig Identities Sin [(pi/20 - x] / Cos [(pi/2) - x]

OpenStudy (john_es):

You should use, \[\sin(a-b)=\sin a\cos b-\cos a \sin b \\ \cos(a-b)=\cos a\cos b+\sin a \sin b\] And, \[\sin (\pi/2)=1\\ \cos(\pi/2)=0\]

OpenStudy (anonymous):

use Cos [(pi/2) - x] = sin x

OpenStudy (anonymous):

how would you use the \[\sin (a - b) and \cos (a - b)\]

OpenStudy (john_es):

Let a=pi/2 and b=x, then, \[\sin(\pi/2-x)=\sin(\pi/2)\cos x-\cos(\pi/2)\sin x\] Now, use \[\sin(\pi/2)=1\\ \cos(\pi/2)=0\] Then, \[\sin(\pi/2-x)=\cos x\] The same follows for the other expression.

OpenStudy (mathmale):

Another possible approach: Note that \[Sin [(\pi/20 - x] / Cos [(\pi/2) - x]\] could be re-written as \[\tan(\frac{ \pi }{ 2 }-x)\]... provided that your pi/20 is a typo and was meant to be pi/2.

OpenStudy (mathmale):

Can this be reduced? Why or why not?